# Global well-posedness of non-integrable hyperbolic-ellptic Ishimori system in the critical Sobolev space

### v1

@@ 4. Estimates for Ishimori equation
If $k\leq k^{\prime}$, we use the bilinear estimate (3.69) to obtain
@@
$\quad\quad+2^{-\sigma k}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{3}\geq k}2^{-\frac{|k_{3}-k^{\prime}|}{2}}b_{k_{3}}b_{k_{3}}(\sigma)$
@@
which yields the desired bound (4.2) in this case.
We further decompose the product $P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])$ according to Bony’s paraproduct structure. For the Low-High interaction $k_{1}\leq k-4$, $|k_{2}-k|\leq 8$ , we have
@@
$\lesssim 2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1}\leq 4}2^{-\frac{|k_{1}-k|}{2}}b_{k_{1}}\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w).$
For the High-Low and High-High interactions, we place the high-frequency term in $L^{2}$ and use the pointwise bound for $P_{k^{\prime}}w$ to obtain
@@
$\sup_{y}\|P_{k}w^{y}P_{k}\mathcal{N}_{1,m}\|$
$\lesssim\sum_{k^{\prime},k_{3}}\int_{0}^{\infty}\sup_{y,y^{\prime}}\|P_{k}w^{y}\cdot P_{k}(\partial_{x}L(P_{k^{\prime}}AP_{k_{3}}\bm{\psi}),L(P_{k^{\prime}}A\partial_{x}P_{k_{3}}\bm{\psi}))\|_{L^{1}_{t,x}}\mathrm{d}s$
@@
$\quad\cdot\int_{0}^{\infty}(2^{2k_{2}})^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}\mathrm{d}s\cdot 2^{-\sigma k_{\max}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)\mathcal{D}(P_{k}w)$
$\lesssim\sum_{k_{i}}2^{\max\{k,k_{3}\}+k_{2}-2\max\{k_{1},k_{2}\}}\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}$
@@
$C(k,k_{1},k_{2},k_{3})\lesssim 2^{-\frac{|\Delta k|}{4}},\ \Delta k:=\max\{k,k_{1},k_{2},k_{3}\}-\min\{k,k_{1},k_{2},k_{3}\},$
@@
If $\max\{k,k_{3}\}\geq\max\{k_{1},k_{2}\}+10$, then the factor vanishes unless $k_{\max}=k_{3}$ and $|k_{3}-k|\leq 10$. In this scenario we have
(4.17) $C(k,k_{1},k_{2},k_{3})\lesssim 2^{k-\max\{k_{1},k_{2}\}}\cdot 2^{-\frac{1}{2}(|k-k_{1}|+|k-k_{2}|)}\lesssim 2^{-\frac{1}{2}|k_{1}-k_{2}|}.$
@@
(4.15) $\lesssim 2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{\max\{k,k_{3}\}\leq\max\{k_{1},k_{2}\}}2^{-\frac{|\Delta k|}{4}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)$
@@
(4.19) $b_{k}(\sigma)+\mathcal{D}(P_{k}\bm{\psi})\lesssim c_{k}(\sigma).$
@@
By the definition (1.46) of $b_{k}(\sigma)$, it suffices to prove the following two estimates:
(4.20) $\mathcal{D}(P_{k}\bm{\psi})\lesssim 2^{-\sigma k}c_{k}(\sigma),$
@@
We apply the Strichartz estimate (2.11) for the ultrahyperbolic Schrödinger equation to the differentiated Ishimori equation (1.19). This yields
@@
It remains to prove the bound for $\psi_{t}$. Usinng the equation (1.17), we have
@@
where $R$ is the Riesz operators. The term $\partial_{x}\bm{\psi}$ is already controlled by the estimate for $\bm{\psi}$. The remaining terms involve products of the connection coefficient $\bm{A}$ and the field $\bm{\psi}$. To estimate these products, we use the following bound for $\bm{A}$, obtained by taking the limit $s\to 0$ in the heat flow estimate (3.42) and using the previously established fact that $a_{k}(\sigma)\lesssim b_{k}(\sigma)\lesssim\varepsilon^{\frac{1}{2}}$:
@@
Applying Lemma 3.1 to the product together with the bound (4.26) and the bootstrap bound for $\bm{\psi}$, we obtain
(4.27) $\|P_{k}(\bm{A}\cdot\bm{\psi},(R^{2}\bm{A})\cdot\bm{\psi})\|_{G}\lesssim 2^{-(\sigma-1)k}c_{k}(\sigma),\ \sigma\in[\frac{1}{5},\sigma_{1}],$
which completes the proof. ∎
@@ Acknowledgements
Y. Zhou was partially supported by NSFC (No. 12171097).

### v2

@@ 4. Estimates for Ishimori equation
If $k\leq k^{\prime}$, for the Low-High and High-Low interactions. where $\max\{k_{1},k_{2}\}\leq k+4$, and for the High-High interactions in the regime $|k_{2}-k^{\prime}|\geq|k-k^{\prime}|$, we use the bilinear estimate (3.69) to obtain
@@
$\quad\quad+2^{-\sigma k}\mathcal{D}(P_{k^{\prime}}w)\sum_{|k_{3}-k^{\prime}|\geq|k-k^{\prime}|}2^{-\frac{|k_{3}-k^{\prime}|}{2}}b_{k_{3}}b_{k_{3}}(\sigma)$
@@
while for the High-High interactions with $|k_{1}-k_{2}|\leq 4,k_{1}\geq k-4$ in the regime $|k_{2}-k^{\prime}|\geq|k-k^{\prime}|$, we note that since $k^{\prime}\geq k$, we have $k_{2}\geq\frac{k+k^{\prime}}{2}$. Placing the low-frequency term in $L^{2}_{t}L^{\infty}_{x}$, we have
$\lesssim\sum_{k_{1},k_{2}}\int_{0}^{\infty}\|P_{k^{\prime}}w\|_{L^{\infty}_{t}L^{2}_{x}}\|P_{k}(P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)])\|_{L^{2}_{t}L^{\infty}_{x}}\mathrm{d}s$
$\lesssim\sum_{k_{1},k_{2}}2^{k}\|P_{k_{1}}\bm{\psi}(s)P_{k_{2}}[D_{x}\bm{\psi}(s)]\|_{L^{2}_{t,x}}\mathcal{D}(P_{k^{\prime}}w)$
$\lesssim\sum_{k_{3}\geq\frac{1}{2}(k+k^{\prime})}2^{k}\mathcal{D}(P_{k^{\prime}}w)\int_{0}^{\infty}(2^{2k_{3}}s)^{-\frac{1}{4}}\langle 2^{2k_{3}}s\rangle^{-6}2^{-\sigma k_{3}}b_{k_{3}}b_{k_{3}}(\sigma)\mathrm{d}s$
$\lesssim 2^{-\sigma k}\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{3}\geq\frac{1}{2}(k+k^{\prime})}2^{k-k_{3}}b_{k_{3}}b_{k_{3}}(\sigma)$
Combing the above subcases yields the desired bound (4.2) when $k^{\prime}\geq k$.
If $k^{\prime}\leq k$, for the Low-High interaction with $k_{1}\leq k-4$, $|k_{2}-k|\leq 8$ , we have
@@
$\lesssim 2^{-\sigma k}2^{-\frac{|k-k^{\prime}|}{2}}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w)\sum_{k_{1}\leq 4}2^{-\frac{|k_{1}-k|}{2}}b_{k_{1}}\lesssim\varepsilon^{\frac{1}{2}}2^{-\sigma k}2^{-\frac{|k-k^{\prime}|}{2}}b_{k}(\sigma)\mathcal{D}(P_{k^{\prime}}w).$
For the High-Low and High-High interactions, we place the high-frequency term in $L_{t,x}^{2}$ and use the pointwise bound for $P_{k^{\prime}}w$ to obtain
@@
$\sup_{y}\|P_{k}w^{y}P_{k}\mathcal{N}_{1,m}\|_{L^{1}_{t,x}}$
$\lesssim\sum_{k^{\prime},k_{3}}\int_{0}^{\infty}\sup_{y}\|P_{k}w^{y}\cdot P_{k}(\partial_{x}L(P_{k^{\prime}}AP_{k_{3}}\bm{\psi}),L(P_{k^{\prime}}A\partial_{x}P_{k_{3}}\bm{\psi}))\|_{L^{1}_{t,x}}\mathrm{d}s$
@@
$\quad\cdot\int_{0}^{\infty}(2^{2k_{2}}s)^{-\frac{1}{4}}\langle 2^{2k_{1}}s\rangle^{-3}\langle 2^{2k_{2}}s\rangle^{-3}\mathrm{d}s\cdot 2^{-\sigma k_{\max}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)\mathcal{D}(P_{k}w)$
$\lesssim\sum_{k_{i}}2^{\max\{k,k_{3}\}+\frac{1}{2}k_{2}-\frac{3}{2}\max\{k_{1},k_{2}\}}\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}$
@@
$C(k,k_{1},k_{2},k_{3})\lesssim\min\{2^{-\frac{1}{2}(|k-k_{1}|+|k_{2}-k_{3}|)},2^{-\frac{1}{2}(|k-k_{2}|+|k_{1}-k_{3}|)}\}$
$\lesssim 2^{-\frac{1}{4}(|k-k_{1}|+|k_{2}-k_{3}|+|k-k_{2}|+|k_{1}-k_{3}|)}$
(4.18) $\lesssim 2^{-\frac{|\Delta k|}{8}},\ \Delta k:=\max\{k,k_{1},k_{2},k_{3}\}-\min\{k,k_{1},k_{2},k_{3}\},$
@@
If $\max\{k,k_{3}\}\geq\max\{k_{1},k_{2}\}+10$, then the factor vanishes unless $k_{\max}=k_{3}$ and $|k_{3}-k|\leq 10$. In this scenario we have $\max\{k_{1},k_{2}\}\leq k+10$ and
(4.19) $C(k,k_{1},k_{2},k_{3})\lesssim 2^{k-\max\{k_{1},k_{2}\}}\cdot 2^{-\frac{1}{2}(|k-k_{1}|+|k-k_{2}|)}$
(4.20) $\lesssim 2^{k-\max\{k_{1},k_{2}\}}\cdot 2^{\frac{1}{2}(k_{1}+k_{2}-2k)}\lesssim 2^{-\frac{1}{2}|k_{1}-k_{2}|}.$
@@
(4.16) $\lesssim 2^{-\sigma k}\mathcal{D}(P_{k}w)\sum_{\max\{k,k_{3}\}\leq\max\{k_{1},k_{2}\}}2^{-\frac{|\Delta k|}{8}}b_{k_{\min}}b_{k_{\text{med}}}b_{k_{\max}}(\sigma)$
@@
(4.22) $b_{k}(\sigma)+\mathcal{D}(P_{k}\bm{\psi})\lesssim c_{k}(\sigma),\ \|P_{k}\psi_{t}\|_{G}\lesssim 2^{k}\varepsilon.$
@@
By the definition (1.47) of $b_{k}(\sigma)$, it suffices to prove the following estimates:
(4.23) $\mathcal{D}(P_{k}\bm{\psi})\lesssim 2^{-\sigma k}c_{k}(\sigma),\ \sigma\in[0,\sigma_{1}]$
@@
(4.25) $\|P_{k}\psi_{t}\|_{G}\lesssim 2^{k}\varepsilon.$
We first prove (4.23). Applying the energy and Strichartz estimate (2.11) for the hyperbolic Schrödinger equation to the differentiated Ishimori equation (1.20), we obtain
@@
It remains to prove (4.24) and (4.25) for $\psi_{t}$. Using the equation (1.18), we have
@@
where $R$ is the Riesz operators. The term $\partial_{x}\bm{\psi}$ is already controlled by the estimate for $\bm{\psi}$. The remaining terms involve products of the connection coefficient $\bm{A}$ and the field $\bm{\psi}$. To estimate these products, we use the following bound for $\bm{A}$, obtained by sending $s\to 0$ in the heat flow estimate (3.42) and using the previously established fact that $a_{k}(\sigma)\lesssim b_{k}(\sigma)\lesssim\varepsilon^{\frac{1}{2}}$:
@@
Applying Lemma 3.1 to the product terms, combined with (4.30) and the bootstrap bound for $\bm{\psi}$, we find for $\sigma\in[\tfrac{1}{5},\sigma_{1}]$ that
$\|P_{k}(\bm{A}\cdot\bm{\psi},(R^{2}\bm{A})\cdot\bm{\psi})\|_{G}\lesssim\sum_{k^{\prime}\leq k}2^{k^{\prime}-\sigma k}\varepsilon^{\frac{1}{2}}b_{k^{\prime}}b_{k}(\sigma)+\sum_{k^{\prime}\geq k}2^{k-\sigma k^{\prime}}\varepsilon^{\frac{1}{2}}b_{k^{\prime}}b_{k^{\prime}}(\sigma)$
(4.31) $\lesssim 2^{-(\sigma-1)k}c_{k}(\sigma)\Big(1+\sum_{k^{\prime}\geq k}2^{\sigma(k-k^{\prime})}2^{\delta(k^{\prime}-k)}\Big)\lesssim 2^{-(\sigma-1)k}c_{k}(\sigma).$
Similarly,
(4.32) $\|P_{k}(\bm{A}\cdot\bm{\psi},(R^{2}\bm{A})\cdot\bm{\psi})\|_{G}\lesssim\sum_{k^{\prime}\leq k}2^{k^{\prime}}\varepsilon^{\frac{1}{2}}b_{k^{\prime}}b_{k}+\sum_{k^{\prime}\geq k}2^{k}\varepsilon^{\frac{1}{2}}b_{k^{\prime}}^{2}\lesssim 2^{k}\varepsilon.$
These estimates control the remaining terms and complete the proof. ∎
@@ Acknowledgements
Y. Zhou was supported by NSFC (No. 12571231, No. 12171097).